# QCResUNet: Joint Subject-level and Voxel-level Segmentation Quality Prediction

### v1

@@ 4.4 Explainability analysis
We hypothesized that the performance gain of the proposed QCResUnet compared to the baselines of UNet, ResNet-34, and ResNet-50 is due to joint training under both subject-level and voxel-level supervision. One potential explanation could be that the joint training allowed the proposed QCResUNet to effectively localize on segmentation errors, as suggested by the attention maps produced by Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2017) (see Figure 7). We observed that the CAM obtained from ResNet-34 and ResNet-50 did not focus on the areas of segmentation errors. Although ResNet-50 had more parameters and depth, it did not demonstrate significant improvement over ResNet-34, indicating that these factors did not improve the QC performance. In contrast, the UNet CAM showed a coarse localization on the difference between the ground truth and the predicted segmentation. In contrast to the baseline methods, the QCResUNet CAM was well localized on segmentation errors.
@@ 5 Discussion and conclusions
Importantly, the proposed method exhibited high generalizability when applied to unseen data, surpassing other state-of-the-art segmentation QC methods. This was particularly true for the RCA and UE-based methods in the brain tumor segmentation QC task, which exhibited poor performance when assessing the quality of segmentation results obtained using segmentation methods different than the ones used to generate training data. The poor generalizability of the RCA method was mainly due to the inherent difficulty of obtaining a representative reference dataset for brain tumor segmentation, which is subject to significant variability. Such significant variability may violate the underlying assumption of the RCA method that there is at least one sample in the reference dataset that can be successfully segmented given a query image-segmentation pair (Robinson et al., 2017; Valindria et al., 2017), which is only valid when dealing with healthy anatomies. In the case of the cardiac segmentation QC task, which involves healthy anatomy, RCA methods achieved better performance compared to the brain tumor case when implemented with the atlas-based segmentation method. Similar to RCA, the UE-based QC did not perform well on subject-level quality prediction as well as localizing segmentation error in the brain tumor case. This may be attributed to the fact that there is inherent variability in uncertainty maps produced by various segmentation methods on datasets with different image quality, tumor characteristics, etc. While in the cardiac segmentation QC task, which involves healthy anatomy and less variability, the UE-based showed better performance. Additionally, as consistent with findings in Jungo et al. (2020), we found that the UE-based method offers limited segmentation error localization. This limitation further hinders its ability to generalize effectively. Furthermore, the UE-based method can only be used to assess segmentations obtained from deep learning models. Unless the deep learning segmentation method directly outputs an estimate of voxel-wise uncertainty, test time estimation of uncertainty (e.g., using MCDropout (Gal and Ghahramani, 2016)) requires access to the model architecture and weights, which may not be possible for models deployed in clinical settings.
@@
The proposed work is not without limitations. Firstly, though the proposed method has better generalizability than other segmentation QC methods, it may be affected by domain shift issues, which are common to all deep learning methods. As a consequence, translating the proposed to clinical practice will require monitoring of each performance to ensure reliable results. In addition, techniques that can enhance the robustness and generalizability of the proposed method to domain shifts (Ganin et al., 2016; Carlucci et al., 2019) are worth further investigating in future work. Second, a key requirement of the proposed method is the availability of all four modalities. Although handling missing modalities enhances the applicability of the proposed method in diverse clinical settings, it is a non-trivial problem. The integration of existing techniques for handling missing modalities (Dorent et al., 2019; Shen and Gao, 2019; Wang et al., 2023) into the proposed QCResUNet will increase significantly the computational burden for training. This is due to the fact that these methods necessitate separate sets of encoders and decoders for each input modality, resulting in a computational load that is at least quadruple that of the current QCResUNet. In our future work, we will explore efficient strategies to allow the model to handle missing modalities. Third, the proposed method was only validated on segmentation tasks involving a single object. Although we have shown that the proposed method generalized well across different segmentation tasks that include multiple tissues, more experiments are needed to evaluate how this method performs in the presence of multiple objects (e.g., multiple lesions). Lastly, this work focused on predicting DSC and NSD as metrics to appropriately summarize segmentation quality. However, it is important to recognize that these metrics have limitations and may not be suitable for all applications. While we demonstrated the proposed method’s ability to handle multiple metrics, additional research will be required to tailor the method for predicting the most appropriate metric for specific tasks (Maier-Hein et al., 2024).
@@ Appendix B Uncertainty estimation-based QC
The uncertainty estimation based QC was implemented using the Monte Carlo Dropout (MCDropout) approach (Gal and Ghahramani, 2016), given its success and popularity. MCDropout was applied to the main convolution blocks in the nnUNet and DeepMedic as well as to the self-attention block in the nnFormer with a drop rate $p=0.5$, following the protocol in (Jungo et al., 2020). After running the inference of the model $T$ times, the uncertainty map of a segmentation can be computed as
@@ B.1 Uncertainty-Error Overlap
We used the uncertainty-error (UE) overlap proposed by Jungo et al. (2020) to calibrate the uncertainty map. The UE overlap measures the overlap between the uncertainty map (UMap) and the ground-truth segmentation error map (SEMgt) following the DSC formula:
@@
where ||$\cdot$|| ⋅ | denotes the cardinality of the set of voxels of each binary map. To compute the UE overlap, the uncertainty map UMap needs to be thresholded into a binary mask. Following the protocol in Jungo et al. (2020), the optimal threshold value was determined by examining different thresholds in the validation set in increments of 0.05, ranging from 0.05 to 0.95. The optimal threshold values for different datasets are shown in Table 1.
@@ B.2 Subject-level DSC prediction
Following Jungo et al. (2020), the voxel-level uncertainty map was aggregated to subject-level Dice Similarity Coefficient (DSC) and Normalized Surface Dice (NSD) predictions by using a two-step approach. First, 102 radiomics features were automatically extracted from the uncertainty map using the PyRadiomics [^1] https://pyradiomics.readthedocs.io package. Second, we trained a random forest regressor using scikit-learn [^2] https://scikit-learn.org/stable/ package in the training dataset to predict the subject-level DSC and NSD.
@@ Appendix C Hyper-parameter tuning results
The hyperparameter tuning was carried out using the Raytune [^3] https://docs.ray.io/en/latest/index.html package. Specifically, we utilized a Bandit-based approach for efficient resource allocations that allowed us to dedicate more resources to the more promising hyperparameter combinations (Li et al., 2018, 2020). The learning curve for Raytune hyperparameter tuning in the BraTS validation set is shown in Figure S1. The optimal hyperparameter combinations determined by Raytune can be found in Table 2.

### v2

@@ 4.4 Explainability analysis
We hypothesized that the performance gain of the proposed QCResUnet compared to the baselines of UNet, ResNet-34, and ResNet-50 is due to joint training under both subject-level and voxel-level supervision. One potential explanation could be that the joint training allowed the proposed QCResUNet to effectively localize on segmentation errors, as suggested by the attention maps produced by Gradient-weighted Class Activation Mapping [Grad-CAM; 52] (see Figure 7). We observed that the CAM obtained from ResNet-34 and ResNet-50 did not focus on the areas of segmentation errors. Although ResNet-50 had more parameters and depth, it did not demonstrate significant improvement over ResNet-34, indicating that these factors did not improve the QC performance. The CAM from UNet is more sparse than that of our model and often highlights regions far away from the tumor where false positive segmentation errors occur. In contrast, the CAM produced by our model consistently highlighted tumorous areas, where significant tissue misclassifications are present. Nonetheless, the CAM produced by our model was more spatially dispersed and blob-like. The difference in spatial support between the UNet CAM and our model’s CAM can be explained by the fact that for our QCResUNet the CAM from the encoder was used, while the CAM from the last layer was used for UNet.
@@ 5 Discussion and conclusions
Importantly, the proposed method exhibited high generalizability when applied to unseen data, surpassing other state-of-the-art segmentation QC methods. This was particularly true for the RCA and UE-based methods in the brain tumor segmentation QC task, which exhibited poor performance when assessing the quality of segmentation results obtained using segmentation methods different than the ones used to generate training data. The poor generalizability of the RCA method was mainly due to the inherent difficulty of obtaining a representative reference dataset for brain tumor segmentation, which is subject to significant variability. Such significant variability may violate the underlying assumption of the RCA method that there is at least one sample in the reference dataset that can be successfully segmented given a query image-segmentation pair [47, 57], which is only valid when dealing with healthy anatomies. In the case of the cardiac segmentation QC task, which involves healthy anatomy, RCA methods achieved better performance compared to the brain tumor case when implemented with the atlas-based segmentation method. Similar to RCA, the UE-based QC did not perform well on subject-level quality prediction as well as localizing segmentation error in the brain tumor case. This may be attributed to the fact that there is inherent variability in uncertainty maps produced by various segmentation methods on datasets with different image quality, tumor characteristics, etc. While in the cardiac segmentation QC task, which involves healthy anatomy and less variability, the UE-based showed better performance. Additionally, as consistent with findings in [24], we found that the UE-based method offers limited segmentation error localization. This limitation further hinders its ability to generalize effectively. Furthermore, the UE-based method can only be used to assess segmentations obtained from deep learning models. Unless the deep learning segmentation method directly outputs an estimate of voxel-wise uncertainty, test time estimation of uncertainty (e.g., using MCDropout [16]) requires access to the model architecture and weights, which may not be possible for models deployed in clinical settings.
Limitations: The proposed work is not without limitations. First, though the proposed method has better generalizability than other segmentation QC methods, it may be affected by domain shift issues, which are common to all deep learning methods. As a consequence, translating the proposed to clinical practice will require monitoring of each performance to ensure reliable results. In addition, techniques that can enhance the robustness and generalizability of the proposed method to domain shifts [17, 10] are worth further investigating in future work.
Second, a key requirement of the proposed method when applied to the brain tumor segmentation case only is the availability of all four modalities. The absence of certain modalities can cause a dramatic drop in QC performance (see more details illustrated in E). However, these four modalities constitute the standard of care for brain tumor diagnosis and monitoring due to the fact that they provide essential and complementary information [7, 6]. Therefore, the requirement does not pose a significant practical limitation in most clinical settings. However, the problem of missing modalities can be addressed in several ways. A straightforward solution could be achieved by training models that take as input different sets of imaging modalities (i.e., have a separate model for each combination of input modalities). However, this is not a scalable solution, requiring $12^{M}–$ – 1 different networks for $M$ different modalities. Alternative approaches exist that handle missing modalities within a single model [see e.g., 13, 53, 59, 44]. However, these require separate sets of encoders and decoders for each input modality, resulting in a computational load at least quadruple that of QCResUNet. Although handling missing modalities would enhance the applicability of the proposed method across diverse clinical settings, it is a non-trivial problem that warrants independent investigation that is beyond the scope of this paper.
@@
Third, the proposed method was only validated on segmentation tasks involving a single object. Although we have shown that the proposed method generalized well across different segmentation tasks that include multiple tissues, more experiments are needed to evaluate how this method performs in the presence of multiple objects (e.g., multiple lesions). Lastly, this work focused on predicting DSC and NSD as metrics to appropriately summarize segmentation quality. However, it is important to recognize that these metrics have limitations and may not be suitable for all applications. While we demonstrated the proposed method’s ability to handle multiple metrics, additional research will be required to tailor the method for predicting the most appropriate metric for specific tasks [38].
@@ Appendix B Uncertainty estimation-based QC
The uncertainty estimation based QC was implemented using the Monte Carlo Dropout (MCDropout) approach [Gal and Ghahramani, 2016], given its success and popularity. MCDropout was applied to the main convolution blocks in the nnUNet and DeepMedic as well as to the self-attention block in the nnFormer with a drop rate $p=0.5$, following the protocol in [Jungo et al., 2020]. After running the inference of the model $T$ times, the uncertainty map of a segmentation can be computed as
@@ B.1 Uncertainty-Error Overlap
We used the uncertainty-error (UE) overlap proposed by Jungo et al. [2020] to calibrate the uncertainty map. The UE overlap measures the overlap between the uncertainty map (UMap) and the ground-truth segmentation error map (SEMgt) following the DSC formula:
@@
where ||$\cdot$|| ⋅ | denotes the cardinality of the set of voxels of each binary map. To compute the UE overlap, the uncertainty map UMap needs to be thresholded into a binary mask. Following the protocol in Jungo et al. [2020], the optimal threshold value was determined by examining different thresholds in the validation set in increments of 0.05, ranging from 0.05 to 0.95. The optimal threshold values for different datasets are shown in Table 1.
@@ B.2 Subject-level DSC prediction
Following Jungo et al. [2020], the voxel-level uncertainty map was aggregated to subject-level Dice Similarity Coefficient (DSC) and Normalized Surface Dice (NSD) predictions by using a two-step approach. First, 102 radiomics features were automatically extracted from the uncertainty map using the PyRadiomics [^1] https://pyradiomics.readthedocs.io package. Second, we trained a random forest regressor using scikit-learn [^2] https://scikit-learn.org/stable/ package in the training dataset to predict the subject-level DSC and NSD.
@@ Appendix C Hyper-parameter tuning results
The hyperparameter tuning was carried out using the Raytune [^3] https://docs.ray.io/en/latest/index.html package. Specifically, we utilized a Bandit-based approach for efficient resource allocations that allowed us to dedicate more resources to the more promising hyperparameter combinations [Li et al., 2018, 2020]. The learning curve for Raytune hyperparameter tuning in the BraTS validation set is shown in Figure S1. The optimal hyperparameter combinations determined by Raytune can be found in Table 3.
@@ Appendix D Voxel-level multi-class segmentation error mask visualization
Appendix E QC performance with missing modalities
[…]
Modalities BraTS-testing BraTS-SSA WUSM
[…]
0.787 0.811 0.499 0.776 0.760 0.387 0.709 0.706 0.338
0.654 0.618 0.515 0.597 0.583 0.409 0.593 0.626 0.371
0.541 0.493 0.475 0.551 0.533 0.384 0.547 0.477 0.350
0.695 0.581 0.429 0.697 0.635 0.363 0.622 0.538 0.349
0.887 0.916 0.669 0.821 0.856 0.628 0.827 0.757 0.600
0.875 0.875 0.546 0.854 0.836 0.494 0.797 0.709 0.444
0.885 0.872 0.536 0.860 0.809 0.490 0.826 0.726 0.447
0.617 0.695 0.588 0.578 0.603 0.520 0.613 0.650 0.480
0.703 0.735 0.599 0.664 0.653 0.506 0.668 0.699 0.479
0.604 0.608 0.448 0.631 0.606 0.485 0.593 0.514 0.449
0.912 0.943 0.715 0.861 0.891 0.638 0.835 0.770 0.594
0.930 0.948 0.709 0.904 0.892 0.624 0.847 0.758 0.592
0.901 0.887 0.547 0.878 0.836 0.497 0.822 0.712 0.446
0.776 0.789 0.629 0.639 0.708 0.515 0.622 0.719 0.489
Average 0.769 0.769 0.565 0.736 0.728 0.496 0.709 0.669 0.459
We reported the subject-level and voxel-level QC performance under the missing-modality settings in Table 2. Compared to the model that takes as input all modalities (as reported in Table 4), the subject-level QC performance on the BraTS internal testing dataset decreased by an average of 19.6% and 20.56% in terms of Pearson correlation coefficients for NSD and DSC prediction, respectively, across all modality combinations. Similarly, on the BraTS-SSA dataset, the Pearson correlation coefficients for NSD and DSC dropped by an average of 22.9% and 22.9%, and by 24.6% and 26.6% on the WUSM dataset. A similar trend was observed for the voxel-level QC performance: the DSC for SEM across all sub-tissue types dropped by an average of 26.5%, 29.3%, and 32.9% on the BraTS testing, BraTS-SSA, and WUSM datasets, respectively.
Although a significant performance drop was observed in the setting of missing modalities, our QCResUNet showed a reasonable performance when at least two modalities are present, achieving an average subject-level Pearson correlation for NSD and DSC of 0.774 and 0.766, respectively, across all datasets. Likewise, the proposed method can achieve an average voxel-level QC performance of 0.547 in terms of DSC for SEM. However, if only one modality is present, especially when T1c or FLAIR is missing, the proposed method showed a relatively poor performance, with an average subject-level Pearson correlation coefficient of 0.609 and 0.543 for NSD and DSC. This is because the T1 and T2 modalities do not provide sufficient information for delineating the enhancing tumor and edema.
Appendix F Computational Complexity
Table 3: Benchmark of computational complexity of different models.
Model Number of Parameters (M) FLOPs (G)
ResNet-34 8.36 249.64
ResNet-50 15.94 302.86
UNet 16.60 670.24
QCResUNet 16.59 296.64
We provided a benchmark of computational complexity for the different architectures we used in this paper in Table 3. The proposed QCResUNet has a parameter and FLOPs count comparable to that of ResNet-50, but nearly doubles the number of parameters of ResNet-34 as a result of the added decoder modules. However, our QCResUNet incorporated a relatively shallow decoder with only three upsampling stages—compared to five in the UNet—resulting in a modest increase in FLOPs relative to ResNet-34, but significantly fewer FLOPs than UNet. We would also like to point out that the proposed attention mechanism is efficient in terms of parameters, as it only involves $1\times 1\times 1$ convolutions.